OpenStudy (anonymous):

BRAIN TEASER: My grandson is about as many days as my son in weeks, and my grandson is as many months as I am in years. My grandson, my son and I together are 120 years. Can you tell me my age in years?

hartnn (hartnn):

x+12x+12*4x =120

hartnn (hartnn):

grandson = 1.96 ?? approx 2 yers...

OpenStudy (anonymous):

g = grandson's age s = son's age m = my age : "My grandson is about as many days as my son is weeks," s = 7g : my grandson is as many months as I am in years. m = 12g : g + s + m = 120 Substitute for m and s g + 7g + 12g = 120 20g = 120 g = 120%2F20 g = 6 yrs is grandson's age then s = 7(6) = 42 yrs is son's age and m = 12(6) = 72 yrs is my age ; : See if we can prove this: "My grandson is about as many days as my son is weeks," a g: 6*365 = 2190 days s: 42*52 = 2184 days, that's why they said "about" : "my grandson is as many months as I am in years." g: 12*6 = 72 months

OpenStudy (anonymous):

Wow, copy and paste. LOL

hartnn (hartnn):

lol, still she got a medal! xP

OpenStudy (anonymous):

By design.

OpenStudy (anonymous):

It's not 72 months. the question says she has a grandson and ur suppose to guess the old person's age!!!

OpenStudy (anonymous):

it's 72 YEARS!!!!